OpenStudy (anonymous):

Can someone explain how to solve for x? I know it'll be put into a quadratic equation and x will have two solutions. I just don't know how to set it up. \frac{2x^2}{2x^2 + 24x} = frac{2}{x}

OpenStudy (anonymous):

Uh. It didn't work. \[\frac{2x^2}{2x^2 + 24x} = \frac{2}{x}\]

OpenStudy (anonymous):

@SWAG @Luis_Rivera @ganeshie8

OpenStudy (anonymous):

I did that and got \[2x^3 = 4x^2 + 48x\]

OpenStudy (anonymous):

Clearly the above isn't a quadratic. I'm so confused.

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):



OpenStudy (anonymous):

Wait, so my xs canceled out? Can you show me using the Equation editor? I understand better that way.

OpenStudy (amistre64):

\[\frac{2x^2}{2x^2 + 24x} = \frac{2}{x}\] \[2x^2(x) = 2(2x^2 + 24x)\] \[2x^3 = 4x^2 + 48x\] \[2x^3 - 4x^2 - 48x=0\]factor our a 2x \[2x(x^2 - 2x - 24)=0\]

OpenStudy (amistre64):

there are at most 3 solutions to the setup that would have to be verified afterwards

OpenStudy (amistre64):

essentially: 2x = 0, when x = 0 x^2 -2x -24 = 0, when x = ... use the quadratic formula as instructed

OpenStudy (anonymous):

Oh, okay, thanks:)

OpenStudy (anonymous):

So I got it down to 4x( x^2 - 4x + 4) However, I ran into a problem. When I plug it into the quadratic equation none of my answers match the answers on the test.